OpenStudy (anonymous):

A piece of rope 5m long is cut into two pieces, on 3/5 as long as the other. How long are the pieces? (I found one but I cant seem to get the other)

OpenStudy (anonymous):

I meant one is 3/5 as long as the other*

OpenStudy (stormfire1):

(3/5)x + 1x = 5 (3/5)x + (5/5)x = 5 (8/5)x = 5 Solve for x to get = 25/8 (or 3 1/8)

OpenStudy (anonymous):

thank you :D